# Large-Area Fabrication-aware Computational Diffractive Optics

### v1

@@ 5.2. Computational Holographic Display
Computational holographic display (CHD) (Chakravarthula et al., 2019; Peng et al., 2020) employs algorithms to simulate optical hologram creation and reconstruction. The computed holographic image is brought to life through a display, usually featuring an illuminating source and a phase-modifying element (DOEs). We focus on 2-D near-field computational holography. Specifically, we use Adam optimizer to solve the optimization problems (4) and (5), where the loss function is defined in (6), together with the training recipe in Section 3.2 (Toy Example). Three 3.6 mm2 DOEs with feature size are designed and fabricated to generate the same 2-D 3.6 mm2 hologram image in 1 cm away from the DOE, for which we build the experimental setup as shown in Fig. 8. The optics in our setup are designed to produce a smooth, spatially clean, and evenly collimated beam (520.6-nm wavelength). This is achieved by using a spatial filter as an intermediate step between expander stage. The resulting beam enables plane-wave illumination for holograms of various sizes. Our physical realization does not require any additional intermediate image plane (such as those implemented with a 4- system to filter out undesirable diffraction orders (Gopakumar et al., 2021)), thus directly assessing the patterns diffracted from DOEs. Finally, for accurate sensor positioning, we employed a motorized linear stage with a positioning precision of . We evaluate 3.6-mm2 DOEs with three different design approaches:
@@ 5.3. Beam Shaping with Diffractive Beam Splitter
Beam shaping, as one of the classic use cases of DOE (Kress and Meyrueis, 2000) has found downstream applications in laser material processing (Kuang et al., 2013), fast LiDAR 3-D sensing (Yuan et al., 2021) and visual vibrometry (Zhang et al., 2023b). Here, we design DOEs as diffractive beam splitters whose goal is to split the collimated laser beam into a regular grid of beams that yields an array of spots with flat-top intensity profiles at the desired plane. This task can be viewed as a special case of computer-generated hologram—instead of generating an image, we steer the DOE to produce an array of focal spots in the destination plane. We find that beam splitter designs without considering the fabrication deformation inevitably lose diffraction efficiency and the spot array intensity uniformity. Again, we design two 3.6 mm2 DOEs with feature size, steering the incident plane-wave beam into a 88 array of flat-top equal-intensity spots in a focal plane located at 1 cm away from the DOE, by minimizing the loss function in (6). We reuse the experimental setup of the holographic display (Fig. 8) and measure the raw intensity patterns for fabricated beam-splitting DOEs under the same lighting condition for fair comparison of diffraction efficiency. The experimental measurements of diffracted spot-array patterns are exhibited in Fig. 12, where the overall spot intensity of the fabrication-aware design is 53% higher than the conventional one, indicating large diffraction efficiency gains brought by our method.
@@ 5.4. Single-DOE Broadband Color Imaging
Broadband color imaging with diffractive optics systems is fundamentally challenging due to strong wavelength-dependent dispersion (Peng et al., 2016; Aieta et al., 2015). High-fidelity broadband reconstruction demands point-spread functions (PSFs) that are simultaneously invertible and spectrally consistent across densely sampled bands (Peng et al., 2016; Chakravarthula et al., 2023; Fröch et al., 2025), which notably increases the memory requirements of the design optimization. In addition, discrepancies between simulated and fabricated PSFs further degrade image quality (Shi et al., 2024b). We validate that the distributed fabrication-aware method addresses both issues.
@@
We also remark that the image quality of our neural designs is already competitive with existing single-DOE computational imaging systems, despite the basic single-step reconstruction method employed here. We find that lightweight, edge-friendly reconstruction methods can be used when an accurate optical design process is taken into account, enabled by our fabrication-aware pipeline. Additional experimental results can be found in Supplementary Materials (Section C.2).
@@ 6. Discussions
We note that the randomness of the fabrication and measurement processes inherently limits our neural lithography model’s precision (forward predictability). To analyze this limitation for our specific process, in preparation for the calibration patterns used to train the neural lithography model (Section 3.2), we deliberately place multiple identical patterns in the design layout periphery. Ideally, these patterns post-fabrication would be identical; however, as shown in Fig. 16 for two patterns, slight differences due to random fabrication and measurement error remain, setting the upper bound of prediction accuracy of the neural lithography model.
@@ 7. Conclusion
We believe our contributions offer a step toward bridging the gap between simulation and mass-market implementation of differentiable imaging optics. Moving forward, incorporating additional fabrication variables, such as material-specific variations, direct-write laser beam, or exploring fabrication-in-the-loop strategies, would be interesting areas for research. We believe these advancements provide a foundation for broader adoption of learned optical systems in real-world imaging and display applications.

### v2

@@ 5.2. Computational Holographic Display
Computational holographic display (CHD) (Chakravarthula et al., 2019; Peng et al., 2020) employs algorithms to simulate optical hologram creation and reconstruction. The computed holographic image is brought to life through a display, usually featuring an illuminating source and a phase-modifying element (DOEs). We focus on 2-D near-field computational holography. Specifically, we use Adam optimizer to solve the optimization problems (4) and (5), where the loss function is defined in (6), together with the training recipe in Section 3.2 (Toy Example). Three 3.63.6 mm2 DOEs with feature size are designed and fabricated to generate the same 2-D 3.63.6 mm2 hologram image in 1 cm away from the DOE, for which we build the experimental setup as shown in Fig. 8. The optics in our setup are designed to produce a smooth, spatially clean, and evenly collimated beam (520.6-nm wavelength). This is achieved by using a spatial filter as an intermediate step between expander stage. The resulting beam enables plane-wave illumination for holograms of various sizes. Our physical realization does not require any additional intermediate image plane (such as those implemented with a 4- system to filter out undesirable diffraction orders (Gopakumar et al., 2021)), thus directly assessing the patterns diffracted from DOEs. Finally, for accurate sensor positioning, we employed a motorized linear stage with a positioning precision of . We evaluate 3.63.6 mm2 DOEs with three different design approaches666Extra results of Xu et al. (2025) and our fabrication-aware approach with 2 neural upsampling are provided in the Supplementary Material.:
@@ 5.3. Beam Shaping with Diffractive Beam Splitter
Beam shaping, as one of the classic use cases of DOE (Kress and Meyrueis, 2000) has found downstream applications in laser material processing (Kuang et al., 2013), fast LiDAR 3-D sensing (Yuan et al., 2021) and visual vibrometry (Zhang et al., 2023b). Here, we design DOEs as diffractive beam splitters whose goal is to split the collimated laser beam into a regular grid of beams that yields an array of spots with flat-top intensity profiles at the desired plane. This task can be viewed as a special case of computer-generated hologram—instead of generating an image, we steer the DOE to produce an array of focal spots in the destination plane. We find that beam splitter designs without considering the fabrication deformation inevitably lose diffraction efficiency and the spot array intensity uniformity. Again, we design two 3.63.6 mm2 DOEs with feature size, steering the incident plane-wave beam into a 88 array of flat-top equal-intensity spots in a focal plane located at 1 cm away from the DOE, by minimizing the loss function in (6). We reuse the experimental setup of the holographic display (Fig. 8) and measure the raw intensity patterns for fabricated beam-splitting DOEs under the same lighting condition for fair comparison of diffraction efficiency. The experimental measurements of diffracted spot-array patterns are exhibited in Fig. 12, where the overall spot intensity of the fabrication-aware design is 53% higher than the conventional one, indicating large diffraction efficiency gains brought by our method.
@@ 5.4. Single-DOE Broadband Color Imaging
Broadband color imaging with diffractive optics systems is fundamentally challenging due to strong wavelength-dependent dispersion (Peng et al., 2016; Aieta et al., 2015). High-fidelity broadband reconstruction demands point-spread functions (PSFs) that are simultaneously invertible and spectrally consistent across densely sampled bands (Peng et al., 2016; Sun et al., 2025; Fröch et al., 2025), which notably increases the memory requirements of the design optimization. In addition, discrepancies between simulated and fabricated PSFs further degrade image quality (Shi et al., 2024b; Chakravarthula et al., 2023). We validate that the distributed fabrication-aware method addresses both issues.
@@
We also remark that the image quality of our neural designs is already competitive with existing single-DOE computational imaging systems, despite the basic single-step reconstruction method with simulated PSFs employed here. We find that lightweight, edge-friendly reconstruction methods can be used when an accurate optical design process is taken into account, enabled by our fabrication-aware pipeline. Additional experimental results, including the comparisons of the inverse-filtering results using simulated and experimental PSFs, can be found in Supplementary Materials (Section C.2).
@@ 6. Discussions
We note that the randomness of the fabrication and measurement processes inherently limits our neural lithography model’s precision (forward predictability). To analyze this limitation for our specific process, in preparation for the calibration patterns used to train the neural lithography model (Section 3.2), we deliberately place multiple identical patterns in the design layout periphery. Ideally, these patterns post-fabrication would be identical; however, as shown in Fig. 16 for two patterns, slight differences due to random fabrication and measurement error remain, setting the upper bound of prediction accuracy of the neural lithography model. Nevertheless, a significant portion of the observed random variation stems from our current academic fabrication facility, which relies on manual interventions such as the control of development time—introducing operator-dependent variability. In commercial foundries, such processes are fully automated with substantially reduced operational variance and higher repeatability.
Model Generalizability
The proprietary and diverse nature of nano/micro-fabrication recipes makes it infeasible to directly apply a model trained in one foundry to another, or even within the same foundry using different materials. Nevertheless, our methodology is universal and can be adapted to other lithography processes through appropriate calibration, rendering it highly suitable for clean-room training. Currently, AFM measurements represent the primary bottleneck in the calibration workflow; however, this requirement remains manageable, as the process can be completed within one day using only 10 relatively small patches.
Camera-in-the-loop (CITL) Approaches
Unlike CITL approaches that learn a global mapping for wave propagation, our digital twin of the manufacturing is designed to learn local shape deviations between the design and the fabricated device. These local variations correspond to process-induced effects such as optical blur, nonlinear material response, and 3D chemical reactions. Our model does not account for global effects. Based on our extensive experience with DOE fabrication, we find such global variations—for instance, those related to position on the wafer—are minimal. This sets nano-fabrication apart from digital holography works employing spatial light modulators (SLM) that may suffer voltage gradients across the chip, and other non-uniformities. The focus on locality offers practical advantages: it allows the model to be trained with limited data and enables rapid recalibration for new material systems or process conditions. In contrast, a full CITL pipeline (for static DOE) with potentially hundreds of prototypes would be intractable in practice, as are AFM measurements for large area samples required to train global models.
@@ 7. Conclusion
We believe our contributions offer a step toward bridging the gap between simulation and mass-market implementation of differentiable imaging optics. Moving forward, incorporating additional fabrication variables, such as material-specific variations, direct-write laser beam, would be interesting areas for research. We believe these advancements provide a foundation for broader adoption of learned optical systems in real-world imaging and display applications.
Acknowledgements.
We thank Dongyu Du and Arturo Burguete Lopez for their supports. Wolfgang Heidrich acknowledges support from KAUST Individual Baseline Funding. Felix Heide was supported by an NSF CAREER Award (2047359), a Packard Foundation Fellowship, a Sloan Research Fellowship, a Disney Research Award, a Sony Young Faculty Award, a Project X Innovation Award, a Bosch Research Award and an Amazon Science Research Award. Fabrication was carried out in the Nanofabrication CoreLabs (NCL) at KAUST.Author contributions:
K.W., W.H. and F.H. conceived the ideas; K.W. designed, implemented and analyzed the system and framework; H.J. and K.W. built the physical setup and captured the experimental results; H.A. established the fabrication pipeline and produced all designed devices; J.S. performed the image reconstruction for the imaging task; K.W., W.H. and F.H. led the manuscript writing; H.J., H.A., J.S., Q.F. assisted in design and analysis, experiments, and manuscript writing; W.H. and F.H. supervised the project.